Find the roots of the equation:

(5.1) z4 + 4 = 0 and z4 − 4 = 0

(5.2) Additional Exercises for practice are given below.

Find the roots of

(a) z8 − 16 = 0

(b) z8 + 16 = 0. 


1
Expert's answer
2021-07-09T08:23:14-0400

Given equation is "z^4 + 4 = 0"

"z^4 = -4"

since, "e^{i\pi} = -1" and let "z = re^{i\theta}"

Then, "(re^{i\theta})^4 = 4e^{i\pi}"

"r^4 e^{i4\theta}= 4e^{i\pi}"

solving it, "r^4 = 4 \implies r=\sqrt{2}"

"4\theta = \pi + 2n\pi \implies \theta = \frac{\pi}{4} + \frac{2n\pi}{4}" where n=0,1,2,3

So, roots of the equation will be,

"z = \sqrt{2}e^{i(\frac{\pi}{4} )}, \sqrt{2}e^{i(\frac{\pi}{4} + \frac{\pi}{2})},\sqrt{2}e^{i(\frac{\pi}{4} + {\pi})},\sqrt{2}e^{i(\frac{\pi}{4} + \frac{3\pi}{2})}"

"z = \sqrt{2} e^{i(\frac{\pi}{4} )},\sqrt{2}e^{i(\frac{3\pi}{4} )},\sqrt{2}e^{i(\frac{5\pi}{4})},\sqrt{2}e^{i(\frac{7\pi}{4} )}"


Since, "e^{i\theta} = cos\theta + isin\theta"

"z_1 =\sqrt{2} e^{i(\frac{\pi}{4} )} = \sqrt{2}(cos\frac{\pi}{4} + isin\frac{\pi}{4}) = \sqrt{2}(\frac{1}{\sqrt{2} } +i\frac{1}{\sqrt{2}}) = 1+i"

"z_2 =\sqrt{2} e^{i(\frac{3\pi}{4} )} = \sqrt{2}(cos\frac{3\pi}{4} + isin\frac{3\pi}{4}) = \sqrt{2}(-\frac{1}{\sqrt{2} } +i\frac{1}{\sqrt{2}}) = -1+i"

"z_3 =\sqrt{2} e^{i(\frac{5\pi}{4} )} = \sqrt{2}(cos\frac{5\pi}{4} + isin\frac{5\pi}{4}) = \sqrt{2}(-\frac{1}{\sqrt{2} } -i\frac{1}{\sqrt{2}}) = -1-i"

"z_4 =\sqrt{2} e^{i(\frac{7\pi}{4} )} = \sqrt{2}(cos\frac{7\pi}{4} + isin\frac{7\pi}{4}) = \sqrt{2}(\frac{1}{\sqrt{2} } -i\frac{1}{\sqrt{2}}) = 1-i"



For "z^4-4 = 0 \implies (z^2-2)(z^2+2)=0"

"z^2-2=0 \implies z = \pm\sqrt{2}"

"z^2+2=0 \implies z^2 = - 2 \implies z=\pm\sqrt{2}i"

Roots are, "z = \sqrt{2},-\sqrt{2},\sqrt{2}i,-\sqrt{2}i"



For "z^8 - 16 = 0 \implies (z^4-4)(z^4+4) = 0"

roots for both are obtained above, so the roots will be,

"z = -\sqrt{2},\sqrt{2},-\sqrt{2}i,\sqrt{2}i,1+i,-1+i,-1-i,1-i"


For "z^8 + 16 = 0"

"z^8 = -16"

"(re^{i\theta})^8 = (16e^{i\pi})"

Solving it, "r = (16)^{1/8} \implies r=\sqrt{2}"

"8\theta = \pi + 2n\pi \implies \theta = \frac{\pi}{8} + \frac{2n\pi}{8 }" where =0,1,2,3,4,5,6,7


"z_1 = \sqrt{2}e^{i\frac{\pi}{8}} = \sqrt{2}(cos\frac{\pi}{8}+isin\frac{\pi}{8})"

"z_2 = \sqrt{2}e^{i\frac{3\pi}{8}} = \sqrt{2}(cos\frac{3\pi}{8}+isin\frac{3\pi}{8})"

"z_3 = \sqrt{2}e^{i\frac{5\pi}{8}} = \sqrt{2}(cos\frac{5\pi}{8}+isin\frac{5\pi}{8})"

"z_4= \sqrt{2}e^{i\frac{7\pi}{8}} = \sqrt{2}(cos\frac{7\pi}{8}+isin\frac{7\pi}{8})"

"z_5 = \sqrt{2}e^{i\frac{9\pi}{8}} = \sqrt{2}(cos\frac{9\pi}{8}+isin\frac{9\pi}{8})"

"z_6 = \sqrt{2}e^{i\frac{11\pi}{8}} = \sqrt{2}(cos\frac{11\pi}{8}+isin\frac{11\pi}{8})"

"z_7 = \sqrt{2}e^{i\frac{13\pi}{8}} = \sqrt{2}(cos\frac{13\pi}{8}+isin\frac{13\pi}{8})"

"z_8 = \sqrt{2}e^{i\frac{15\pi}{8}} = \sqrt{2}(cos\frac{15\pi}{8}+isin\frac{15\pi}{8})"